OpenStudy (anonymous):

Factoring the equation will give the binomial: Select one: a. (x – 5)(x + 5) b. (x + 5)(x +10) c. (x – 5)(x – 10) d. None of the above

OpenStudy (anonymous):

\[\sqrt{x-1}=x-7\]

OpenStudy (precal):

square both sides you will get x-1=(x-7)^2 now expand it